OpenStudy (anonymous):

Which is the correct step below for using the grouping method to factor the following polynomial? 5xy + 15x - 2y - 6 A. x(5y + 15) - 1(2y + 6) B. 5x(y + 3) + 2(y + 3) C. 5x(y + 3) - 2(y + 3) D. 5x(y - 3) - 2(y - 3)

OpenStudy (anonymous):

I think its B but not sure?

OpenStudy (whpalmer4):

Well, expand B, what do you get? \[5x(y+3) + 2(y+3) =\]

OpenStudy (whpalmer4):

Hint: for it to be a correct step, it must not change the value of the thing being factored!

OpenStudy (whpalmer4):

Another hint, which you probably don't need, is that the two groups, after factoring, must have a common factor (so A is not an option)

OpenStudy (anonymous):

So I was right? sorry it took me so long to reply

OpenStudy (whpalmer4):

Expand B. What do you get?

OpenStudy (anonymous):

5x(y + 3) + 2(y + 3) this is what I got. Is that what you wanted? Sorry I just learned how to do this! :P

OpenStudy (whpalmer4):

Multiply it out!

OpenStudy (anonymous):

(5x+2)(y+3)?

OpenStudy (whpalmer4):

NO... What is \[5x(y+3)=\] What is \2(y+3)=\] Add them together. What do you get?

OpenStudy (whpalmer4):

Sorry, second equation should be \[2(y+3)=\]

OpenStudy (anonymous):

5xy+15x+2y+6

OpenStudy (whpalmer4):

Yes, and is that what you set out to factor?

OpenStudy (anonymous):

ohhh I see! I added them and I got what the original polynomial was!

OpenStudy (anonymous):

yes!

OpenStudy (whpalmer4):

No, it's not! \[5xy + 15x - 2y - 6\] is what you are supposed to factor!

OpenStudy (anonymous):

oh...... hehe! im really bad at math if you haven't noticed already.

OpenStudy (whpalmer4):

bad at math may not really be true so much as a failure to look very closely...

OpenStudy (anonymous):

yeah your right!

OpenStudy (whpalmer4):

convince yourself of the correctness of each step along the way and you'll magically get a lot better at math :-)

OpenStudy (whpalmer4):

so, what's your new choice for the correct answer? check it in the same fashion as we did here...

OpenStudy (anonymous):

I am thinking C?

OpenStudy (whpalmer4):

If you checked it in the same way, you should be more certain than that :-)

OpenStudy (whpalmer4):

\[5x(y + 3) - 2(y + 3) = 5x*y + 5x*3 -2*y -2*3 = 5xy+15x-2y-6\] That looks like a good answer!

OpenStudy (anonymous):

I have a 99.8% that its C!

OpenStudy (whpalmer4):

There you go. An old math teacher of mine liked to respond to answers such as "I am thinking C?" with "is that an answer, or a prayer?" :-)

OpenStudy (anonymous):

So it is 100% C!?!

OpenStudy (whpalmer4):

Very good! So the next step here would be \[5x(y+3)-2(y+3) = (y+3)(5x-2)\] And we check our work: \[(y+3)(5x-2) = y(5x-2) + 3(5x-2) = 5xy -2y + 15x -6 = 5xy + 15x -2y -6\checkmark\]

OpenStudy (whpalmer4):

@#$@#% OS. \[5xy-2y+15x-6 = 5xy+15x-2y-6\checkmark\]

OpenStudy (whpalmer4):

Do you have any questions about this factoring by grouping business?

OpenStudy (anonymous):

Thank you soooooo much! Nope! Thank you! :D

OpenStudy (whpalmer4):

Okay, remember to check your work with a skeptical eye :-)